OpenStudy (anonymous):

Sarah went to an electronics store, purchased 3 CDs and 5 DVDs, and spent $137. Michelle went to the same store, bought 2 CDs and 7 DVDs, and spent $161. How much does each CD sell for at the store? a) $13 b)$14 c)$18 d) $19

OpenStudy (anonymous):

(b)

OpenStudy (anonymous):

Let x = price of CD and y = price of DVD Therefore 3x + 5y = 137 ............(1) for Sarah 2x + 7y = 167 ............(2) for Michelle then u got simultaneous equations Subtract (2) from (1) to get: x - 2y = -24 solve for y: y = 0.5x + 12 ............(3) Substitute (3) into either (1) or (2) and solve for x (price of CD) It should give you $14